# Graph Learning

### v1

@@ 5.1.2 Sequential Generative Graph Learning
One core challenge in sequential generating methods is how to decompose the graph into multiple components to facilitate their sequential generation, i.e., determining the generation order of $s_{i}$. Based on the type of generation sequence, sequential generating methods can be further divided into four categories. Node-Sequence-Based, Edge-Sequence-Based, Motif-Sequence-Based, and Rule-Sequence-Based.
@@
Node-Sequence-Based
@@
(1) Collective-Associated-Edge-Based chen2021order predict the entire edge set $\{e_{i,j}\}_{j<i}$ in one operation by considering the relationships between node $v_{i}$ and existing nodes in $G_{i}$. Most studies rely on the adjacency vector $\textbf{a}_{i}$ to predict all potential edges between the newly generated node $v_{i}$ and the existing nodes. Therefore, each generative unit can be further represented as $s_{i}=(v_{i},\textbf{a}_{i})$, and the sequence can be expressed as $Seq(G)=\{(v_{1},\textbf{a}_{1}),\dots,(v_{N},\textbf{a}_{N})\}$.
@@
(2) Progressive-Associated-Edge-Based : The time complexity of the collective-associated-edge-Based method is $O(N^{2})$, which can be computationally expensive for sparse graphs. To address this, progressive-associated-edge generation [liu2018constrained] has been proposed. This method incrementally selects nodes from the existing node set $v_{<i}$ to connect with the current node $v_{i}$, continuing until the desired number of connections is reached. Specifically, it first determines the size of the edge set $\{e_{i,j}\}_{j<i}$ for node $v_{i}$ and then selects nodes to connect from $v_{<i}$. If $M$ represents the number of edges, then the time complexity for this method is denoted by $O(MN)$.
@@
Edge-Sequence-Based
@@
Motif-Sequence-Based
@@
(1) Domain-Agnostic-Based These methods [liao2019efficient, jang2023hierarchical] aim to generate general graphs without requiring domain-specific knowledge. Similar to collective-associated-edge-Based methods in node-sequence-based generation, these methods generate an adjacency vector for each edge. However, the key difference is that while collective-associated-edge-Based methods produce a single node at each step, domain-agnostic-based methods generate multiple nodes at each step.
(2) Domain-Specific-Based Domain-Specific-Based methods [jin2020hierarchical, liu2023using] involve using domain knowledge to define graph motifs and their connections, such as molecular fragments [jin2018junction] or structural motifs [jin2020hierarchical] in molecular graphs, or temporal motifs [liu2023using] in temporal networks. These motifs help define how to connect the generated motifs to construct a complete graph.
@@
Rule-Sequence-Based
@@
dai2018syntax introduces a semantic constraint component during the parse tree generation phase, enabling the generator to ensure both syntactic and semantic validity. wang2024grammar introduce Grammar Prompting, which allows LLMs to use domain-specific constraints and external knowledge during in-context learning. Grammar Prompting enhances model capabilities by incorporating a specialized grammar into each example demonstration.
@@ 5.2.1 Conditional Graphs Generative Graph Learning
For the retrosynthesis prediction task, shi2020graph propose the G2Gs model, which splits a molecule into synthons by breaking the reaction center. Then, a series of graph transformations is applied to generate the corresponding reactants. zhong2023retrosynthesis propose the Graph2Edits model for retrosynthesis prediction, aimed at performing molecular graph editing. wu2021graph formalize the handwritten mathematical expression recognition problem as a graph-to-graph learning task and proposed a G2G model, which has shown significant improvement in experimental results.
In other research directions, hou2024dag propose the DAVA model, which generates social propagation graphs by integrating topological analysis of propagation graphs with corresponding user attributes as prior knowledge. guo2020interpretable introduce the NED-VAE model to tackle disentangled representation learning for attributed graphs.
@@ 5.2.3 Conditional Sequence Generative Graph Learning
In the NLP domain, conditional sequence generative graph learning is applied to tasks such as Semantic Parsing and mind map generation. bai2022graph consider enhancing AMR parsing and generation by pretraining on semantic graphs. hu2021efficient propose an EMGN that considers document-level semantics through sequence-to-graph modeling. zhang2024coreference introduce a CMGN to incorporate external structural knowledge for mind map generation.
@@
In other directions, conditional sequence generative graph learning methods have been designed to address specific challenges. For instance, in the context of privacy preservation, xiang2023privacy propose a Gromov-Wasserstein Autoregressive (GWAR) model, which does not require correspondence information, thereby maintaining the privacy of nodes in evolving graphs. For synthetic dataset challenges, zheng2024temporal introduce the concept of Time-Bounded Communities to represent communities that are destined to disband. Based on patterns collected from real-world data, they design GTB, a temporal graph Generation method featuring Time-Bound communities.
@@
Providing explanations for graph learning models is an effective manner to enhance trust in their predictions. Consequently, explainable graph learning has received a lot of attention from the graph learning research community in recent years [yucage, huang2025sehg, nandan2025graphxai]. Unlike explaining traditional deep learning models, there are challenges specific to explaining graph learning models. First, the topological information of graph data is complex. Its node features and topological relations impact on the prediction results are more difficult to be captured compared to structured data. Second, the predictions of graph neural networks contain higher-order dependencies. The multi-layer message aggregation mechanism of graph neural networks causes the representation of nodes to depend on the information of multi-hop neighbors, which raises the difficulty of explaining graph neural networks. As a result of these challenges, explainable graph learning has been a key research topic in the graph learning field, and numerous methods have been presented to improve the explainability of graph learning. To this end, we consider explainable graph learning as a branch of graph learning and provide a systematic survey of studies within this branch, thereby providing an updated overview.
@@ 6.1.1 Factual Explanation
Generative-based. Generative-based explainers typically utilize generative models to generate new graph explanation. The generative model learns representations of input graphs and constrains the maximization of mutual information between the graph representations and predictions to fit a probability distribution of the target model. Furthermore, the generative model reconstructs the explainable subgraphs based on the learned graph representations. Numerous existing explainers fall into this type. For instance, lin2021generative utilize the variation in model loss caused by removing edges from the input graph to assess the causal effect of edges on predictions, which in turn guides the generation of explanatory subgraphs. In addition, some studies attempt to optimize graph editing strategies with reinforcement learning for guiding the generation of explanatory subgraphs. yuan2020xgnn utilize reinforcement learning to guide the editing of the input graph (removing or adding edges) for constraining the explainer to generate subgraphs that contain critical information. Similarly, NEURIPS2021_be26abe7 utilize reinforcement learning to guide the decoder for reconstructing explanatory subgraphs, and leverage a fully connected layer to capture the representation of the input graph. Besides reinforcement learning methods, some studies [wang2023gnninterpreter, li2023dag] constrain the explanatory subgraphs to contain critical topological information by learning probabilistic generative graph distributions to approximate the spatial distance between the representation of the explanatory subgraph and the representation of the input graph. Despite the superior generalization and computational efficiency of generative explainers, subgraph generation processes are still black-box since they require learning the graph representation.
@@
Search-based. In contrast to black-box generative explainers, search-based explainers identify the importance of edges and nodes in the input graph through a white-box search strategy without learning the representation of the input graph. For instance, some studies [ying2019gnnexplainer, luo2020parameterized, wang2021towards, funke2022zorro, yuan2021explainability, zhang2022gstarx] inject perturbations to each edge and node on the input graph for searching subgraph structures that significantly influence the prediction. The explainer measures the influence on predictions by perturbing different edges and nodes to capture important topological information. Inspired by the change rate of gradient representations, pope2019explainability measure the sensitivity of predictions to the change rate of gradient representations from topological information. Theoretically, predictions are sensitive to gradient changes in key topological information. Therefore, the explainer quantifies such sensitivity to capture the key explanatory subgraphs. In addition, some explainers [feng2022degree, xiong2023relevant] decompose the predictions of the target model and back-propagate the decomposition values layer by layer up to the input layer. The input layer searches for a feature subset that critically influences the prediction according to the decomposition values.
@@ 6.1.2 Counterfactual Explanation
Learning-based. Learning-based methods employ generative models to capture the feature distribution of input graphs and reconstruct counterfactual graphs that closely resemble the original. Given their ability to handle complex graph data, these methods offer excellent scalability and generalization. Consequently, numerous learning-based counterfactual explainers have been proposed. For instance, bajaj2021robust propose an unsupervised learning method to capture the decision boundaries of each category, thereby fitting a common decision logic for the target model across categories. By learning this decision logic, the explainer identifies subsets of edges that influence the prediction. These edge subsets are then removed from the input graph to generate counterfactual explanations. ma2022clear utilize a graph variational auto-encoder to encode the input graph representation and reconstruct the graph representation into a counterfactual explainable graph.
@@ 6.2.1 Information Extraction
As shown in Figure 6.5, information extraction approaches typically use mutual information as a constraint to guide subgraph extraction. Specifically, self-explanatory methods based on information extraction aim to maximize the mutual information between subgraphs and the final prediction while minimizing it between subgraphs and the input graphs. These methods encourage subgraphs to discard irrelevant features from the original graph, retaining only those key features that significantly impact prediction. For example, DBLP:conf/icml/MiaoLL22 propose a method namely GSAT (Graph Stochastic Attention) that uses a stochastic attention mechanism to calculate the variational upper bound for $I(G,G_{s})$, where $G_{s}$ is an informative subgraph from the input graph $G$. Similar to GSAT, LRI (Learnable Randomness Injection)  [DBLP:conf/iclr/0001LLL23] uses both Bernoulli and Gaussian distribution as the prior distribution. LRI separately calculates the existence and location importance of points to study how perturbing the location of the point will affect the predictions. DBLP:conf/iclr/YuXRBHH21 propose a bi-level optimization method, which uses the Donsker-Vardhan KL representation to estimate $I(G,G_{s})$ in the inner loop and optimize the GIB (Graph Information Bottleneck) objective in the outer loop. To improve the efficiency and stability of this bi-level training process, DBLP:conf/cvpr/Yu0H22 propose another method named VGIB (Variational Graph Information Bottleneck). VGIB uses a different compression technique by injecting noise into the node representations via a learned probability for each node, and the higher classification loss stands for the informative substructure injected with noise. Thus, the informative substructure is less likely to be injected with noise compared to label-irrelevant substructures. 10.1145/3589334.3645674 propose a self-explanatory model based on graph distribution matching, which reconstructs a new graph by matching the distributions of the original graph and the explanatory subgraph in its feature space during model training to capture the most informative subgraph. NEURIPS2023_f224f056 propose a prototype-based graph information bottleneck that uses prototype learning to capture key subgraphs in the input graph, which significantly influence the prediction, thereby enabling the prototypes to encapsulate richer key substructures.
@@ 6.2.2 Structure Extraction
Unlike information extraction, structure extraction focuses on constraining the topology of the graph instead of mutual information. As shown in Figure 6.6, structure extraction-based self-explanatory approaches directly constrain the topology in the input graph to identify key subgraphs that maximize the accuracy of the predictions made by the graph learning model. For example, DBLP:conf/iclr/WuWZ0C22 propose to generate explanations by dividing the input into two parts: a causal component $(C)$ and a non-causal component $(S)$. The method is trained to minimize classification loss while ensuring that the prediction $Y$ is conditionally independent of $S$ given $C$. This is accomplished by constructing several interventional distributions during training, where the features that remain stable across these interventions are identified as causal. The process begins by selecting the top-$k$ edges from the input graph to form the causal portion, leaving the rest as non-causal. A distribution intervention module then perturbs the data to help uncover invariant causal patterns. Lastly, two classifiers are applied to produce a combined prediction based on both parts. DBLP:conf/aaai/ZhangLWLL22 propose a method, namely ProtoGNN (Prototype Graph Neural Network), that combines prototype learning [DBLP:journals/ijmi/SchmidtMBPG01] with GNNs. Instead of relying on post-hoc explanation techniques, ProtoGNN makes predictions by comparing input graphs to a set of learned prototypes in the latent space, and allows explanations to emerge directly from the case-based reasoning process. DBLP:conf/cikm/DaiW21 introduce a method that identifies the $k$ nearest labeled nodes based on both structural and feature similarities. These nearest neighbors are leveraged not only for making predictions but also for producing explanations. An explanation subgraph is constructed by selecting nodes whose importance exceeds a predefined threshold. DBLP:conf/aaai/FengY0T22 propose KER-GNN (Kernel Graph Neural Networks), a framework that incorporates graph kernels into the GNN message-passing mechanism. In this approach, node representations are updated by measuring the similarity between each node’s local subgraph and a set of trainable hidden graph filters. These learnable filters capture key structural patterns within the data, and the resulting node embeddings can be further used to identify significant substructures for interpretation.
@@
Fairness in graph learning refers to the elimination of any bias or favouritism toward specific individuals or groups in model predictions [dong2023fairness, luo2025fairgp]. As an extension of deep neural networks to the graph domain, GNNs are typically data-driven and aimed at learning mappings from graph-based data. However, the learning process, particularly mechanisms like message-passing, can inadvertently reinforce existing biases in the data, potentially leading to unfair predictions influenced by sensitive attributes such as race or gender [li2024rethinking]. Therefore, how to tackle unfairness issues in graph learning algorithms naturally becomes another crucial responsible problem. In this survey, we review responsible graph learning with a particular focus on Privacy-Preserving Graph Learning and Fairness in Graph Learning.

### v2

@@ 5.1.2 Sequential Generative Graph Learning
One core challenge in sequential generation methods is how to decompose the graph into multiple components to facilitate their sequential generation, i.e., determining the generation order of $s_{i}$. Based on the type of generation sequence, sequential generation methods can be further divided into four categories, that are node-Sequence-Based methods, edge-Sequence-based methods, motif-Sequence-based methods, and rule-sequence-based methods.
@@
Node-Sequence-Based Methods.
@@
(1) Collective-Associated-Edge-Based Methods. chen2021order predicted the entire edge set $\{e_{i,j}\}_{j<i}$ in one operation by considering the relationships between node $v_{i}$ and existing nodes in $G_{i}$. Most studies rely on the adjacency vector $\textbf{a}_{i}$ to predict all potential edges between the newly generated node $v_{i}$ and the existing nodes. Therefore, each generative unit can be further represented as $s_{i}=(v_{i},\textbf{a}_{i})$, and the sequence can be expressed as $Seq(G)=\{(v_{1},\textbf{a}_{1}),\dots,(v_{N},\textbf{a}_{N})\}$.
@@
(2) Progressive-Associated-Edge-Based Methods. The time complexity of the collective-associated-edge-Based method is $O(N^{2})$, which can be computationally expensive for sparse graphs. To address this, progressive-associated-edge generation [liu2018constrained] has been proposed. This method incrementally selects nodes from the existing node set $v_{<i}$ to connect with the current node $v_{i}$, continuing until the desired number of connections is reached. Specifically, it first determines the size of the edge set $\{e_{i,j}\}_{j<i}$ for node $v_{i}$ and then selects nodes to connect from $v_{<i}$. If $M$ represents the number of edges, then the time complexity for this method is denoted by $O(MN)$.
@@
Edge-Sequence-Based Methods.
@@
Motif-Sequence-Based Methods.
@@
(1) Domain-Agnostic-Based Methods. These methods [liao2019efficient, jang2023hierarchical] aim to generate general graphs without requiring domain-specific knowledge. Similar to collective-associated-edge-Based methods in node-sequence-based generation, these methods generate an adjacency vector for each edge. However, the key difference is that while collective-associated-edge-Based methods produce a single node at each step, domain-agnostic-based methods generate multiple nodes at each step.
(2) Domain-Specific-Based Methods. Domain-Specific-Based methods [jin2020hierarchical, liu2023using] involve using domain knowledge to define graph motifs and their connections, such as molecular fragments [jin2018junction] or structural motifs [jin2020hierarchical] in molecular graphs, or temporal motifs [liu2023using] in temporal networks. These motifs help define how to connect the generated motifs to construct a complete graph.
@@
Rule-Sequence-Based Methods.
@@
dai2018syntax introduces a semantic constraint component during the parse tree generation phase, enabling the generator to ensure both syntactic and semantic validity. wang2024grammar introduced Grammar Prompting, which allows LLMs to use domain-specific constraints and external knowledge during in-context learning. Grammar Prompting enhances model capabilities by incorporating a specialized grammar into each example demonstration.
@@ 5.2.1 Conditional Graphs Generative Graph Learning
For the retrosynthesis prediction task, shi2020graph proposed the G2Gs model, which splits a molecule into synthons by breaking the reaction center. Then, a series of graph transformations is applied to generate the corresponding reactants. zhong2023retrosynthesis proposed the Graph2Edits model for retrosynthesis prediction, aimed at performing molecular graph editing. wu2021graph formalized the handwritten mathematical expression recognition problem as a graph-to-graph learning task and proposed a G2G model, which has shown significant improvement in experimental results.
In other research directions, hou2024dag proposed the DAVA model, which generates social propagation graphs by integrating topological analysis of propagation graphs with corresponding user attributes as prior knowledge. guo2020interpretable introduced the NED-VAE model to tackle disentangled representation learning for attributed graphs.
@@ 5.2.3 Conditional Sequence Generative Graph Learning
In the NLP domain, conditional sequence generative graph learning is applied to tasks such as Semantic Parsing and mind map generation. bai2022graph considered enhancing AMR parsing and generation by pretraining on semantic graphs. hu2021efficient proposed an EMGN that considers document-level semantics through sequence-to-graph modeling. zhang2024coreference introduced a CMGN to incorporate external structural knowledge for mind map generation.
@@
In other directions, conditional sequence generative graph learning methods have been designed to address specific challenges. For instance, in the context of privacy preservation, xiang2023privacy proposed a Gromov-Wasserstein Autoregressive (GWAR) model, which does not require correspondence information, thereby maintaining the privacy of nodes in evolving graphs. For synthetic dataset challenges, zheng2024temporal introduced the concept of Time-Bounded Communities to represent communities that are destined to disband. Based on patterns collected from real-world data, they design GTB, a temporal graph Generation method featuring Time-Bound communities.
@@
Providing explanations for graph learning models is an effective manner to enhance trust in their predictions. Consequently, explainable graph learning has received a lot of attention from the graph learning research community in recent years [yucage, huang2025sehg, nandan2025graphxai] and has been applied in various fields. Unlike explaining traditional deep learning models, there are challenges specific to explaining graph learning models. First, the topological information of graph data is complex. Its node features and topological relations impact on the prediction results are more difficult to be captured compared to structured data. Second, the predictions of graph neural networks contain higher-order dependencies. The multi-layer message aggregation mechanism of graph neural networks causes the representation of nodes to depend on the information of multi-hop neighbors, which raises the difficulty of explaining graph neural networks. As a result of these challenges, explainable graph learning has been a key research topic in the graph learning field, and numerous methods have been presented to improve the explainability of graph learning. To this end, we consider explainable graph learning as a branch of graph learning and provide a systematic survey of studies within this branch, thereby providing an updated overview.
@@ 6.1.1 Factual Explanation
Generative-based. Generative-based explainers typically utilize generative models to generate new graph explanations. The generative model learns representations of input graphs and constrains the maximization of mutual information between the graph representations and predictions to fit a probability distribution of the target model. Furthermore, the generative model reconstructs the explainable subgraphs based on the learned graph representations. Numerous existing explainers fall into this type. For instance, lin2021generative utilized the variation in model loss caused by removing edges from the input graph to assess the causal effect of edges on predictions, which in turn guides the generation of explanatory subgraphs. In addition, some studies attempt to optimize graph editing strategies with reinforcement learning for guiding the generation of explanatory subgraphs. yuan2020xgnn utilized reinforcement learning to guide the editing of the input graph (removing or adding edges) for constraining the explainer to generate subgraphs that contain critical information. Similarly, NEURIPS2021_be26abe7 utilized reinforcement learning to guide the decoder for reconstructing explanatory subgraphs, and leverage a fully connected layer to capture the representation of the input graph. Besides reinforcement learning methods, some studies [wang2023gnninterpreter, li2023dag] constrain the explanatory subgraphs to contain critical topological information by learning probabilistic generative graph distributions to approximate the spatial distance between the representation of the explanatory subgraph and the representation of the input graph. Despite the superior generalization and computational efficiency of generative explainers, subgraph generation processes are still black-box since they require learning the graph representation.
@@
Search-based. In contrast to black-box generative explainers, search-based explainers identify the importance of edges and nodes in the input graph through a white-box search strategy without learning the representation of the input graph. For instance, some studies [ying2019gnnexplainer, luo2020parameterized, wang2021towards, funke2022zorro, yuan2021explainability, zhang2022gstarx] inject perturbations to each edge and node on the input graph for searching subgraph structures that significantly influence the prediction. The explainer measures the influence on predictions by perturbing different edges and nodes to capture important topological information. Inspired by the change rate of gradient representations, pope2019explainability measured the sensitivity of predictions to the change rate of gradient representations from topological information. Theoretically, predictions are sensitive to gradient changes in key topological information. Therefore, the explainer quantifies such sensitivity to capture the key explanatory subgraphs. In addition, some explainers [feng2022degree, xiong2023relevant] decompose the predictions of the target model and back-propagate the decomposition values layer by layer up to the input layer. The input layer searches for a feature subset that critically influences the prediction according to the decomposition values.
@@ 6.1.2 Counterfactual Explanation
Learning-based. Learning-based methods employ generative models to capture the feature distribution of input graphs and reconstruct counterfactual graphs that closely resemble the original. Given their ability to handle complex graph data, these methods offer excellent scalability and generalization. Consequently, numerous learning-based counterfactual explainers have been proposed. For instance, bajaj2021robust proposed an unsupervised learning method to capture the decision boundaries of each category, thereby fitting a common decision logic for the target model across categories. By learning this decision logic, the explainer identifies subsets of edges that influence the prediction. These edge subsets are then removed from the input graph to generate counterfactual explanations. ma2022clear utilized a graph variational auto-encoder to encode the input graph representation and reconstruct the graph representation into a counterfactual explainable graph.
@@ 6.2.1 Information Extraction
As shown in Figure 6.5, information extraction approaches typically use mutual information as a constraint to guide subgraph extraction. Specifically, self-explanatory methods based on information extraction aim to maximize the mutual information between subgraphs and the final prediction while minimizing it between subgraphs and the input graphs. These methods encourage subgraphs to discard irrelevant features from the original graph, retaining only those key features that significantly impact prediction. For example, DBLP:conf/icml/MiaoLL22 proposed a method namely GSAT (Graph Stochastic Attention) that uses a stochastic attention mechanism to calculate the variational upper bound for $I(G,G_{s})$, where $G_{s}$ is an informative subgraph from the input graph $G$. Similar to GSAT, LRI (Learnable Randomness Injection) [DBLP:conf/iclr/0001LLL23] uses both Bernoulli and Gaussian distribution as the prior distribution. LRI separately calculates the existence and location importance of points to study how perturbing the location of the point will affect the predictions. DBLP:conf/iclr/YuXRBHH21 proposed a bi-level optimization method, which uses the Donsker-Vardhan KL representation to estimate $I(G,G_{s})$ in the inner loop and optimizes the GIB (Graph Information Bottleneck) objective in the outer loop. To improve the efficiency and stability of this bi-level training process, DBLP:conf/cvpr/Yu0H22 proposed another method named VGIB (Variational Graph Information Bottleneck). VGIB uses a different compression technique by injecting noise into the node representations via a learned probability for each node, and the higher classification loss stands for the informative substructure injected with noise. Thus, the informative substructure is less likely to be injected with noise compared to label-irrelevant substructures. 10.1145/3589334.3645674 proposed a self-explanatory model based on graph distribution matching, which reconstructs a new graph by matching the distributions of the original graph and the explanatory subgraph in its feature space during model training to capture the most informative subgraph. NEURIPS2023_f224f056 proposed a prototype-based graph information bottleneck that uses prototype learning to capture key subgraphs in the input graph, which significantly influence the prediction, thereby enabling the prototypes to encapsulate richer key substructures.
@@ 6.2.2 Structure Extraction
Unlike information extraction, structure extraction focuses on constraining the topology of the graph instead of mutual information. As shown in Figure 6.6, structure extraction-based self-explanatory approaches directly constrain the topology in the input graph to identify key subgraphs that maximize the accuracy of the predictions made by the graph learning model. For example, DBLP:conf/iclr/WuWZ0C22 proposed to generate explanations by dividing the input into two parts: a causal component $(C)$ and a non-causal component $(S)$. The method is trained to minimize classification loss while ensuring that the prediction $Y$ is conditionally independent of $S$ given $C$. This is accomplished by constructing several interventional distributions during training, where the features that remain stable across these interventions are identified as causal. The process begins by selecting the top-$k$ edges from the input graph to form the causal portion, leaving the rest as non-causal. A distribution intervention module then perturbs the data to help uncover invariant causal patterns. Lastly, two classifiers are applied to produce a combined prediction based on both parts. DBLP:conf/aaai/ZhangLWLL22 proposed a method, namely ProtoGNN (Prototype Graph Neural Network), that combines prototype learning [DBLP:journals/ijmi/SchmidtMBPG01] with GNNs. Instead of relying on post-hoc explanation techniques, ProtoGNN makes predictions by comparing input graphs to a set of learned prototypes in the latent space, and allows explanations to emerge directly from the case-based reasoning process. DBLP:conf/cikm/DaiW21 introduced a method that identifies the $k$ nearest labeled nodes based on both structural and feature similarities. These nearest neighbors are leveraged not only for making predictions but also for producing explanations. An explanation subgraph is constructed by selecting nodes whose importance exceeds a predefined threshold. DBLP:conf/aaai/FengY0T22 proposed KER-GNN (Kernel Graph Neural Networks), a framework that incorporates graph kernels into the GNN message-passing mechanism. In this approach, node representations are updated by measuring the similarity between each node’s local subgraph and a set of trainable hidden graph filters. These learnable filters capture key structural patterns within the data, and the resulting node embeddings can be further used to identify significant substructures for interpretation.
@@
As an extension of neural networks on graph-structured data, GNNs are also vulnerable to adversarial attacks [tang2025rmtrans]. Due to their unique message-passing mechanism and graph structure, GNNs can be negatively affected by adversarial perturbations on both graph structures and node attributes. The vulnerability of GNNs has raised tremendous concerns about their adoption in safety-critical domains such as credit estimation and healthcare. For instance, fraudsters could create several transactions with deliberately chosen high-credit users to escape GNN-based fraud detectors, which could cause significant losses to individuals and institutions. Hence, developing robust GNNs is another important aspect of trustworthiness, and many efforts have been taken. This area is mature enough that several comprehensive surveys about adversarial attacks and defenses on graphs already exist. In this section, we provide a brief overview of adversarial learning on graphs but focus more on methods in emerging directions such as scalable attacks, graph backdoor attacks, and recent defense methods.
Fairness in graph learning refers to the elimination of any bias or favouritism toward specific individuals or groups in model predictions [dong2023fairness, luo2025fairgp]. Therefore, both individual fairness and group fairness need to be taken into account. To this end, different types of datasets have been proposed for studying different types of fairness.
As an extension of deep neural networks to the graph domain, GNNs are typically data-driven and aimed at learning mappings from graph-based data. However, the learning process, particularly mechanisms like message-passing, can inadvertently reinforce existing biases in the data, potentially leading to unfair predictions influenced by sensitive attributes such as race or gender [li2024rethinking]. Therefore, how to tackle unfairness issues in graph learning algorithms naturally becomes another crucial responsible problem. In this survey, we review responsible graph learning with a particular focus on Privacy-Preserving Graph Learning and Fairness in Graph Learning.